Five years ago, Anu was thrice as old as Sonu. After ten years, Anu will be twice as old as Sonu. How old are Anu and Sonu?

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Let us assume the present ages of Anu is x and Sonu is y.

According to the question:

x – 5 = 3(y – 5)

x – 5 = 3y – 15

x= 3y-15+5

x= 3y -10……………….. (1)

Again, as per question;

x + 10 = 2(y + 10)

x + 10 = 2y + 20

x – 2y = 10

(3y – 10) – 2y = 10 (from equation 1)

3y-2y =10+10

y= 20

Substitute, y=20 in equation 1,

x = 3(20) -20

x =60 – 20

x = 50

Hence, the present ages of Anu is 50 years and of Sonu is 20 years.


Thanks